OpenStudy (anonymous):

About Multiple Integral integral of (x^2+xy^3)dxdy R=[2,3]*[4,7]

terenzreignz (terenzreignz):

X goes from 2 to 3 and Y goes from 4 to 7?

terenzreignz (terenzreignz):

\[\int\limits_{x = 2}^{3}\int\limits_{y = 4}^{7}(x^{2} + xy^{3})dxdy\] Is this what you meant?

OpenStudy (anonymous):

Yes! you are right. how about the integral?

OpenStudy (unklerhaukus):

\[\int\limits_4^7\int\limits_2^3 (x^2+xy^3) \text dx\text dy\]

terenzreignz (terenzreignz):

Well, the trick here is to just evaluate one integral at a time, and with respect to only one variable, first. In other words, you either integrate as if y is a constant, or as if x is a constant. It's your call Can you do it? :)

OpenStudy (unklerhaukus):

make sure you take the limits in the right order ,

OpenStudy (unklerhaukus):

\[=\int\limits_4^7\left(\int\limits_2^3 (x^2+xy^3) \text dx\right)\text dy\]

OpenStudy (unklerhaukus):

treat y as a constant in the inner integral

OpenStudy (anonymous):

=int[7,4](x^3/3+x^2y^3/2|[3,2])dy =

OpenStudy (unklerhaukus):

good \[=\int\limits_4^7\left(\left.\frac{x^3}{3}+\frac{x^2y^3}{2}\right|_2^3\right)dy\] now take the inner limits for x,